half the sum of two numbers is 17 if the difference between the numbers is 18 whar are the numbers?
Answer by mananth(16949)

half the sum of the numbers = 17
sum of numbers will be 34
let one number be x
the other will be
34-x
x-(34-x) =18
x-34+x= 18
2x= 34+18
2x=52
x= 26
the other number will be 34-26=8